OpenStudy (anonymous):

Simplify the expression. (-2i)(8i)

OpenStudy (anonymous):

-16i^2

OpenStudy (shamil98):

Remember i = √-1 so i^2 = -1 you can simplify further

OpenStudy (shamil98):

It becomes positive 16.

OpenStudy (anonymous):

can you please show me how you got that?

OpenStudy (anonymous):

-1 x -16

OpenStudy (shamil98):

(-2i)(8i) is simply multiplication -2 x 8 = -16 i x i = -1 this becomes -16 x - 1 = 16

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

8i * 2i 8 * 2 * i * i 16 * i^2 i^2 = -1. 16 * -1 = -16